OpenStudy (anonymous):

a=1/2h(B+b) solve for B

OpenStudy (anonymous):

I assume two possibilities: a) That h is part of the numerator A = 1/2 h (B+ b) 2A = h (B+b) 2A / h = B + b 2A/ h - b = B b) That h is part of the denominator A = 1/2 h (B = b) 2Ah = B + b 2Ah - b = B

OpenStudy (anonymous):

please vote for my design http://www.signazon.com/contest/grad2014/lkkajqe5/

OpenStudy (anonymous):

did that hekp

OpenStudy (anonymous):

please vote guys i hlpd u please vote for my design http://www.signazon.com/contest/grad2014/lkkajqe5/